OpenStudy (anonymous):

Solve by addition or substitution y=4x-4y 3x+2y=11 Help.

OpenStudy (anonymous):

multiply 3x+2y=11 all by -2 and line it up with y=4x-4y. then solve for one variable and then use replacement.

OpenStudy (anonymous):

so plug in -2 for x and y and multiply 11 as well by -2?

OpenStudy (anonymous):

everything by -2(...)

OpenStudy (anonymous):

@Ben103 We already went through a similar question like this. This just contains different numbers.

OpenStudy (unklerhaukus):

\[y=4x−4y\qquad\qquad3x+2y=11\] a different method would be to solve the first equation for x (by taking the y terms to the left and dividing by four) then substitute x into the second equation

OpenStudy (anonymous):

So far I have (4x-44)-(-6x+-4x)=(-22)-44. Is that good so far?

OpenStudy (unklerhaukus):

are you sure you have the question right?

OpenStudy (unklerhaukus):

i dont think there are nice solutions to this system

OpenStudy (anonymous):

I multiplied -2 by (3x+2y) and 11 and got -6x+-4x=-22

OpenStudy (anonymous):

and then lined it up with y=4x-4y

OpenStudy (anonymous):

You don't use the elimination method in this one although it still works. YOu use the substitution method because one of the equations has y as the subject which makes your life easier when solving using substitution.